# Various Properties of Various Ultrafilters, Various Graph Width Parameters, and Various Connectivity Systems

### v6

@@ A.11.1 Hierarchy for branch-width
  • For every graph $G$ with $\text{bw}(G)\neq 0$, $\text{boolw}(G)\leq\text{bw}(G)$ (Boolean-width) [5].
@@ A.13 Rank-width
The concept of a rank function is introduced from the world of matroids, and parameters such as Branch-width, Tree-width, and Tangles have been extended from graphs to matroids[330, 577, 304, 394, 395]. Therefore, we will briefly introduce the concept of matroids. For those interested in learning more about matroids, please refer to books, lecture notes, or surveys on the topic[579, 697, 607, 700, 329].
@@
Matroid can also be defined as follows.
@@
Based on the above, we define Rank-width as follows.
@@ A.14.1 Hierarchy for clique-width
  • For any graph $G$, $\text{fun}(G)\leq 2\text{clique-width}(G)-1$. (fun: functionality). [20, 224]
Proof A.31.
@@ A.15 Linear-Clique-width
Linear Clique-width is the linear restriction of Clique-width.
@@ A.18 Hypertree-width
The known obstructions are Hypertangles [12] and Hyperbrambles [12]. In the future, we will consider a hyperultrafilter analogous to [304] (cf.[525]).
@@ A.18.1 Related concepts for Hypertree-width
  • Hyper-T-width and Hyper-D-width [617]: These graph width parameters were introduced as the first stable connectivity measures for hypergraphs, offering tools for designing algorithms to address various problems in hypergraph theory.
@@ A.20 Submodular-width
Submodular-width [344] is defined using the submodular condition property.
@@ A.26 Twin-width
Twin-width [729, 733, 731, 732, 730, 103] is a graph width parameter that measures the complexity of graphs by evaluating their ability to avoid a fixed minor. This includes classes such as planar graphs and graphs with bounded tree-width or clique-width.
@@
Let $G=(V,E,R)$ be a trigraph, where $V$ is the vertex set, $E$ is the set of black edges, and $R$ is the set of red edges. For any two vertices $u,v\in V$, the contraction of $u$ and $v$ in $G$, denoted $G/u,v=(V^{\prime},E^{\prime},R^{\prime})$, is defined as follows:
@@ A.31 Path-distance-width
Path distance width (PDW) [715] measures the width from a distance perspective, constrained by bandwidth. Related concepts include Rooted Path Distance width (RPDW) [715], Connected Path-distance-width, and c-connected Path-distance-width [506, 575]. Distance width is particularly useful in problems like Graph Isomorphism, where the goal is to determine if two graphs are identical by checking if there is a one-to-one correspondence between their vertices and edges that preserves the connections (cf.[606, 355, 499]).
@@ A.32.1 Hierarchy for Tree-partition-width
Proof A.58.
@@
Related concepts are Directed Tree partition width [414], which considers Tree partition width in Digraphs.
@@ A.35 Dominating-set-width
Dominating-set-width [536] measures the number of different dominating sets on a subgraph.
@@ A.36 Point width
Point width [138] for hypergraphs provides a condition for the tractability of Max-CSPs, generalizing bounded MIM-width and $\beta$-acyclicity.
@@ A.40.1 Hierarchy for Directed Tree-width
  • If a digraph $G$ has bounded DAG-width, then it also has bounded Directed Tree-width [62].
@@
  • For every digraph $D$, it holds that $\text{dtw}(D)\geq\text{cycle-degeneracy}(D)$.
Proof A.68.
@@ A.41.1 Hierarchy for Directed Path-width
    Directed Vertex Separation Number: The Directed Path-width is equal to the Directed Vertex Separation Number [82].
@@
  3. 3.
    Entanglement: The entanglement of $D$ is less than or equal to the Directed Path-width [598] (cf.[139]).
@@
  4. 4.
    Cycle Rank: The Directed Path-width is less than or equal to the Cycle Rank of $D$ [357].
@@
  7. 7.
    Diblockage: A directed graph has Directed Path-width $\geq k-1$ if and only if it has a diblockage of order $\geq k$ [243].
@@ A.51.1 Hierarchy for Stretch-width
  • The stretch-width of any graph is at most twice its clique-width.
@@ A.53 Perfect matching width
Perfect matching width [337] is a width parameter for matching covered graphs based on a branch decomposition.
@@ A.54 Bisection-width
The Bisection problem seeks to partition the vertices of a graph into two equally sized sets while minimizing the cut size. The Bisection Width is a width parameter closely related to this Bisection problem [64, 643, 66]. If a graph G has bounded Band-width, then it also has bounded bisection-width [668].
@@ A.58 Query-width
Query-width is graph width parameter of a common generalization of both acyclicity and tree-width [96, 263].
@@ A.59.1 Related concepts for Universal width
  • Split-width [16, 174]: Width parameter for an alternating finite automaton (AFA).
@@ A.63 Cycle-width
Cycle-width [386] measures graph complexity, focusing on cycle structures and highlighting differences between directed and undirected tree-width by emphasizing local graph properties.

### v8

@@ A.11.1 Hierarchy for branch-width
[…]
@@ A.13 Rank-width
The concept of a rank function is introduced from the world of matroids, and parameters such as Branch-width, Tree-width, and Tangles have been extended from graphs to matroids[565, 966, 529, 662, 663]. Therefore, we will briefly introduce the concept of matroids. For those interested in learning more about matroids, please refer to books, lecture notes, or surveys on the topic[968, 1208, 1029, 1213].
@@
Matroid can also be defined as follows[968, 1208].
@@
Based on the above, we define Rank-width as follows [682] .
@@ A.14.1 Hierarchy for clique-width
[…]
  • The stretch-width of any graph is at most twice its clique-width[1255].
  • Fusion-width(G) $\leq$ Clique-width(G) [532].
  • If a graph $G$ has bounded edge clique cover number, then it also has bounded Clique-width (and simultaneous interval number)[116].
@@
  • For any graph $G$, $\text{fun}(G)\leq 2\text{clique-width}(G)-1$. (fun: functionality). [51, 399]  [^37] Functionality of a vertex $y$ is the minimum number of vertices $k$ such that $y$ can be represented as a Boolean function of $k$ vertices[51, 399].
Proof A.36.
@@ A.14.2 Related concepts for clique-width
  • courcelle2004clique: Symmetric clique-width is a graph parameter that measures complexity using equivalence classes of vertex adjacency relationships within subsets.
@@ A.15 Linear-Clique-width
Linear Clique-width (Also called Sequential clique-width[446, 447]) is the linear restriction of Clique-width.
@@ A.18 Hypertree-width
Example A.43.
Let $H$ be a hypergraph with vertex set $V(H)=\{A,B,C,D,E\}$ and hyperedge set $E(H)=\{e_{1},e_{2},e_{3}\}$, where:
$e_{1}=\{A,D\},\quad e_{2}=\{D,E\},\quad e_{3}=\{A,B,C\}.$
Thus, $H$ is represented by the pair $H(V,E)=(\{A,B,C,D,E\},\{\{A,D\},\{D,E\},\{A,B,C\}\})$.
@@
The known obstructions are Hypertangles [16] and Hyperbrambles [16]. In the future, we will consider a hyperultrafilter analogous to [529] (cf.[872, 509]).
@@ A.18.1 Related concepts for Hypertree-width
  • Hyper-T-width and Hyper-D-width [1047]: These graph width parameters were introduced as the first stable connectivity measures for hypergraphs, providing tools for designing algorithms to address various problems in hypergraph theory. They are closely related to Hyper-Directed-width.
@@
  • FAQ-width [2, 759] : FAQ-width is a width parameter used to optimize the evaluation of functional aggregate queries (FAQs) with free variables over semirings, generalizing fractional hypertree width[2, 759].
  • m-width [710] : m-width is a graph width parameter used in join processing, often smaller than fractional hypertree width, aiding in identifying subquadratically solvable joins[710].
@@
  • Biconnected width [480]: Biconnected width measures the maximum vertex count in biconnected components of a hypergraph’s primal graph.
  • TCLUSTER-width [321, 586]: TCLUSTER-width measures the tree-likeness of a hypergraph, defined as the maximum clique size in its chordal graph representation.
  • Hinge-width [624, 625]: Hinge-width measures a hypergraph’s complexity by generalizing acyclic hypergraphs, computable in polynomial time.
  • Hierarchical hypertree width [671]: Hierarchical hypertree width measures how far a join query is from being hierarchical, influencing the complexity of temporal join algorithms [671].
  • Nest set-width [820] : Nest set-width is a graph width parameter defined by the smallest width of a nest set elimination order, which organizes vertices into subsets ordered by inclusion[820].
  • $\beta$-Hyperorder width [587, 242]: $\beta$-Hyperorder width is a graph width parameter that generalizes hypertree width, tailored for analyzing the complexity of negative conjunctive queries [587, 242]. A related parameter is $\beta$-fractional hyperorder width.
  • # hypertree-width [273, 595] : # hypertree-width is a structural measure of query instances that determines the tractability of counting solutions in conjunctive queries and constraint satisfaction problems[273, 595].
  • Primal treewidth [82] : Primal treewidth is the treewidth of the primal graph of a CSP instance, where vertices represent variables and edges connect variables appearing together in constraints[82].
    Related parameters are Incidence tree-width, modular incidence treewidth[559], Incidence clique-width, twinclass-pathwidth[650], twinclass-treewidth[650], modular tree-depth[939], prefix pathwidth[405], (generalized) modular tree-width[802], modular pathwidth[895], and Incidence MIM-width[1126, 1214].
    Various hierarchies are also known. For example, $\beta$-hypertree width is less than or equal to incidence clique-width[559]. Incidence clique-width is less than or equal to signed incidence clique-width[457] and modular incidence treewidth[559]. Signed incidence clique-width[457] and modular incidence treewidth[980, 559] are less than or equal to incidence treewidth[1126, 1054]. Incidence treewidth is less than or equal to primal treewidth[1054].
Question A.45.
Is there any new characteristic that emerges by extending each of these width parameters to SuperHyperGraphs? (cf.[521, 509])
@@ A.20 Submodular-width
Submodular-width [583] is defined using the submodular condition property. And sharp submodular width is a graph width parameter that refines submodular width, allowing improved algorithm runtimes for evaluating queries on arbitrary semirings[758].
@@ A.26 Twin-width
Twin-width [1256, 1260, 1258, 1259, 1257, 187] is a graph width parameter that measures the complexity of graphs by evaluating their ability to avoid a fixed minor. Intuitively, twin-width reflects the extent to which a graph resembles a cograph—a type of graph that can be reduced to a single vertex by repeatedly merging twin vertices, which are vertices sharing the same neighbors. This width includes classes such as planar graphs and graphs with bounded tree-width or clique-width.
@@
(cf.[1256, 1260, 1258, 1259, 1257, 187] ) Let $G=(V,E,R)$ be a trigraph, where $V$ is the vertex set, $E$ is the set of black edges, and $R$ is the set of red edges. For any two vertices $u,v\in V$, the contraction of $u$ and $v$ in $G$, denoted $G/u,v=(V^{\prime},E^{\prime},R^{\prime})$, is defined as follows:
@@ A.31 Path-distance-width
Path distance width (PDW) [1233, 64, 1230] measures the width from a distance perspective, constrained by bandwidth. Related concepts include Rooted Path Distance width (RPDW) [1233], Connected Path-distance-width, and c-connected Path-distance-width [844, 964]. Distance width is particularly useful in problems like Graph Isomorphism, where the goal is to determine if two graphs are identical by checking if there is a one-to-one correspondence between their vertices and edges that preserves the connections (cf.[1028, 601, 831]).
@@ A.32.1 Hierarchy for Tree-partition-width
  • $tn(G)\leq 3\times\text{tpw}(G)$ (tn: track-number) [387].
  • $qn(G)\leq 3\times\text{tpw}(G)-1$ (qn: queue-number) [387].
Proof A.65.
@@
Related concepts are Directed Tree partition width [696], which considers Tree partition width in Digraphs, and star-partition-width [1216]
@@ A.35 Dominating-set-width
Dominating-set-width [899] measures the number of different dominating sets on a subgraph.  [^43] A dominating set in a graph is a subset of vertices such that every vertex in the graph is either in the dominating set or adjacent to a vertex in it [1194].
@@ A.36 Point width
Point width [243] for hypergraphs provides a condition for the tractability of Max-CSPs, generalizing bounded MIM-width and $\beta$-acyclicity.  [^44] In graph algorithms, tractability refers to whether a problem can be solved efficiently, typically in polynomial time. Tractable problems have algorithms that run in reasonable time, while intractable problems are computationally difficult. And Max-CSPs (Maximum Constraint Satisfaction Problems) involve finding an assignment to variables that maximizes the number of satisfied constraints in a given set. These problems are NP-hard and widely studied(cf.[1042]).
@@ A.40 Directed Tree-width
The definition of Directed Tree-width is provided below. Note that various papers have proposed and examined alternative definitions of Directed Tree-width.
@@
A similar concept to tree-width is DAG-width, which we define below [130]. If a graph has entanglement $k$, its DAG-width is at most $k+1$[130].
Definition A.74.
[130] Let $G=(V,E)$ be a graph. A set $W\subseteq V$ guards a set $V^{\prime}\subseteq V$ if, whenever there is an edge $(u,v)\in E$ such that $u\in V^{\prime}$ and $v\notin V^{\prime}$, then $v\in W$.
Definition A.75 (DAG-decomposition).
[130] Let $G=(V,E)$ be a directed graph. A DAG-decomposition is a tuple $D=(D,\{X_{d}\}_{d\in V(D)})$ such that:
  • (D1)
    $D$ is a directed acyclic graph (DAG).
  • (D2)
    $\bigcup_{d\in V(D)}X_{d}=V$.
  • (D3)
    For all $d\preceq d^{\prime}\preceq d^{\prime\prime}$ in $D$, we have $X_{d}\cap X_{d^{\prime\prime}}\subseteq X_{d^{\prime}}$.
  • (D4)
    For a root $d$, $X_{d}$ is guarded by $\emptyset$.
  • (D5)
    For all $(d,d^{\prime})\in E(D)$, $X_{d}\cap X_{d^{\prime}}$ guards $X_{d^{\prime}}\setminus X_{d}$, where $X_{d^{\prime}}:=\bigcup_{d^{\prime}\preceq d^{\prime\prime}}X_{d^{\prime\prime}}$.
The width of a DAG-decomposition $D$ is defined as $\max\{|X_{d}|:d\in V(D)\}$. The DAG-width of a graph is the minimal width over all possible DAG-decompositions of $G$.
@@ A.40.1 Hierarchy for Directed Tree-width
  • If a digraph $G$ has bounded DAG-width, then it also has bounded Directed Tree-width [129]  [^45] DAG-width measures how close a directed graph is to being acyclic. It is defined as the minimal width of a DAG-decomposition, which is similar in concept to tree-decomposition in undirected graphs[129, 957, 130, 647]. DAG-width is also used in Parity game. Related parameters of parity game are alternation-depth, nesting depth [756, 755]..
@@
  • For every digraph $D$, it holds that $\text{dtw}(D)\geq\text{cycle-degeneracy}(D)$. [^46] Cycle-degeneracy (or c-degeneracy) of a directed graph is the minimum number of vertices that must be removed to eliminate all directed cycles containing any given vertex[953].
  • For every digraph $G$, $1/3(dtw(G)+2)\leq\textit{entanglement}(G)$[712].
Proof A.77.
@@ A.41.1 Hierarchy for Directed Path-width
    Directed Vertex Separation Number: The Directed Path-width is equal to the Directed Vertex Separation Number [161].  [^47] The Directed Vertex Separation Number of a digraph measures the smallest number of vertices needed to separate the remaining vertices into disjoint sets, ensuring no directed path crosses between these sets[161, 171].
@@
    Directed feedback vertex set number: If $G$ has bounded Directed feedback vertex set number, then $G$ has bounded Directed Path-width [950].
  3. 3.
@@
  4. 4.
    Entanglement: The entanglement of $D$ is less than or equal to the Directed Path-width [1009] . [^48] Entanglement of a graph measures the complexity of its cycles, defined by the minimum number of detectives required to catch a thief in a graph-based game[247, 132].
@@
  5. 5.
    Cycle Rank: The Directed Path-width is less than or equal to the Cycle Rank of $D$ [604].  [^49] Cycle rank is a measure of the complexity of a connected graph, defined as the number of independent cycles in the graph.
@@
  8. 8.
    Diblockage: A directed graph has Directed Path-width $\geq k-1$ if and only if it has a diblockage of order $\geq k$ [429].  [^50] Diblockage is directed version of blockage.
@@ A.51.1 Hierarchy for Stretch-width
  • The stretch-width of any graph is at most twice its clique-width[1255].
@@ A.53 Perfect matching width
Perfect matching width [574] is a width parameter for matching covered graphs based on a branch decomposition. 515151 A perfect matching in a graph is a set of edges in which every vertex is paired with exactly one other vertex. A matching covered graph is a connected graph where every edge is part of at least one perfect matching, meaning all edges are ”admissible.” These concepts are central to matching theory, a topic extensively studied in graph theory research [1127, 648, 715, 249, 1248].
@@ A.54 Bisection-width
The Bisection problem seeks to partition the vertices of a graph into two equally sized sets while minimizing the cut size. The Bisection Width is a width parameter closely related to this Bisection problem [135, $1$, 137]. If a graph $G$ has bounded Band-width, then it also has bounded bisection-width [1154].
@@ A.58 Query-width
Query-width is a graph width parameter that generalizes both acyclicity and tree-width, measuring the complexity of hypergraphs through restricted hypertree-width[178, 461, 585, 270].
@@ A.59.1 Related concepts for Universal width
  • Split-width [23, 309]: Width parameter for an alternating finite automaton (AFA). Also Split-width is a graph parameter used to measure the complexity of the behavior graphs of computational models, aiding in decidability results for systems like multi-pushdown systems [308, 307, 24].
@@ A.61 Median-width
The definition of a median graph is provided below for reference.
Definition A.105.
A median graph is an undirected graph $G=(V,E)$ such that, for any three vertices $a,b,c\in V$, there exists a unique vertex $m(a,b,c)\in V$ (called the median) that lies on the shortest paths between each pair of the three vertices. That is, $m(a,b,c)$ belongs to the shortest paths between $a$ and $b$, $b$ and $c$, and $a$ and $c$.
Formally, a vertex $m(a,b,c)$ satisfies:
$d_{G}(a,m(a,b,c))+d_{G}(m(a,b,c),b)=d_{G}(a,b),$
$d_{G}(b,m(a,b,c))+d_{G}(m(a,b,c),c)=d_{G}(b,c),$
$d_{G}(a,m(a,b,c))+d_{G}(m(a,b,c),c)=d_{G}(a,c),$
where $d_{G}(x,y)$ denotes the shortest path distance between vertices $x$ and $y$ in $G$.
@@ A.63 Cycle-width
Cycle-width [646, 1203] measures graph complexity, focusing on cycle structures and highlighting differences between directed and undirected tree-width by emphasizing local graph properties.
Question A.106.
Is there a corresponding width parameter for Walk-width and Circuit-width? If not, can it be defined?
A.63.1 Related concepts for cycle-width
Related concepts is following.
  • radial cycle-width [49]: Radial cycle-width measures the smallest radial width in a cycle-decomposition of a graph, where the radial width is the largest radius of the cycle’s parts.
Question A.107.
Is it possible to define Directed Cycle-Width by considering directed cycles?
Question A.108.
Can Cycle-width be extended to Fuzzy Graphs or similar structures?
A.63.2 Hierarchy for Cycle-width
We consider about Hierarchy for Cycle-width.
Theorem A.109.
The relationships between cycle-width and other graph parameters are as follows:
  • The cycle-width of a graph $G$ is at most $k-1$ if and only if its arc thickness is at most $k$ [1203].
  • For any graph $G$, $\frac{1}{2}\cdot\text{pathwidth}(G)\leq\text{cyclewidth}(G)\leq\text{pathwidth}(G)$[1203].
  • Let $T$ be a tree. Then $\text{cyclewidth}(T)=\text{pathwidth}(T)$[1203].
Proof A.110.
Please refer to each reference.
@@ A.64.1 Related concepts for arc-width
Question A.112.
Is it possible to define Directed Arc-Width by considering directed arcs?
Question A.113.
Can Arc-width be extended to Fuzzy Graphs or similar structures?
@@ A.68 Plane-width
A.69 Scan-width
Scan-width[670, 669] is a width parameter for directed acyclic graphs (DAGs) that measures their tree-likeness by considering the direction of arcs(cf.[1069]).
A.70 V-width
V-width is a complexity measure in graph structures that is smaller than, and potentially much smaller than, the maximum clique size and the maximum number of parents per node in a DAG[283].
A.71 Cross-width
Cross-width is a graph parameter used in kernelization, particularly in analyzing the complexity of reducing problems, often related to polynomial equivalence relations[695].
A.72 Guidance-width
Guidance-width measures the minimal number of colors needed to color a guidance system in a tree, ensuring that conflicting guides are distinctly colored [179].
A.73 Star-width
Star-width measures the complexity of a graph’s decomposition into "stars" (graphs with a shared central vertex) where the width is determined by the size of the interfaces connecting the stars[158, 1172, 1171]. Related parameters are radial star-width[49], costar-width[158], and atomic star-width[158]. If a graph has bounded starwidth, it has bounded tree-depth[1171].
Definition A.114.
[1172, 1171] A star decomposition of a graph $G=(V,E)$ is a star with a core $X_{0}$ and leaves $X_{1},\dots,X_{m}$, where each $X_{i}$ is associated with a subset of $V$, satisfying the following properties:
  1. 1.
    The union of all sets $X_{i}$ equals $V$.
  2. 2.
    For every edge $(v,w)\in E$, there exists a subset $X_{i}$ that contains both $v$ and $w$.
  3. 3.
    For $i\neq j$, if $X_{i}$ and $X_{j}$ both contain a vertex $v$, then $X_{0}$ contains $v$ as well.
The width of a star decomposition is defined as $\max_{i}|X_{i}|-1$. The starwidth $sw(G)$ of a graph $G$ is the minimum width among all possible star decompositions of $G$.
Question A.115.
Is it possible to define Directed Star-Width?
Question A.116.
Can Star-width be extended to Fuzzy Graphs or similar structures?
@@ B.1 Tree-length and Branch-length
And similar concepts are tree-stretch and tree-distortion. Also, please provide the statement to be written as a theorem.
Theorem B.4.
[378] Let $G=(V,E)$ be an undirected, unweighted graph with $n$ vertices. The following inequalities hold for various graph parameters:
  1. 1.
    The tree-breadth $\text{tb}(G)$ is bounded by the tree-length $\text{tl}(G)$ as follows:
    $\text{tb}(G)\leq\text{tl}(G)\leq 2\cdot\text{tb}(G).$
    Similarly, the cluster-radius $R_{s}(G)$ and cluster-diameter $\Delta_{s}(G)$ are related as:
    $R_{s}(G)\leq\Delta_{s}(G)\leq 2\cdot R_{s}(G).$
  2. 2.
    The hyperbolicity $\text{hb}(G)$ of the graph is related to the tree-length and cluster-diameter by:
    $\text{hb}(G)\leq\text{tl}(G)\leq O(\text{hb}(G)\cdot\log n),$
    and
    $\text{hb}(G)\leq\Delta_{s}(G)\leq O(\text{hb}(G)\cdot\log n).$
    [280, 281]
  3. 3.
    The tree-stretch $\text{ts}(G)$ and tree-distortion $\text{td}(G)$ are related to the cluster-diameter as follows:
    $\text{ts}(G)\geq\text{td}(G)\geq\frac{1}{3}\Delta_{s}(G),$
    and
    $\text{td}(G)\leq 2\cdot\Delta_{s}(G)+2,\quad\text{for every }s\in V.$
    [282]
  4. 4.
    The cluster-radius $R_{s}(G)$ is bounded by the tree-distortion as follows:
    $R_{s}(G)\leq\max\{3\cdot\text{td}(G)-1,2\cdot\text{td}(G)+1\},\quad\text{for every }s\in V.$
    [282]
  5. 5.
    The tree-length $\text{tl}(G)$, cluster-diameter $\Delta_{s}(G)$, and cluster-radius $R_{s}(G)$ are related by:
    $\text{tl}(G)-1\leq\Delta_{s}(G)\leq 3\cdot\text{tl}(G),$
    and
    $R_{s}(G)\leq 2\cdot\text{tl}(G),\quad\text{for every }s\in V.$
    [370, 369]
  6. 6.
    The tree-breadth $\text{tb}(G)$ is related to the cluster-radius as follows:
    $\text{tb}(G)-1\leq R_{s}(G)\leq 3\cdot\text{tb}(G).$
    [379]
  7. 7.
    The tree-length $\text{tl}(G)$, tree-distortion $\text{td}(G)$, and tree-stretch $\text{ts}(G)$ satisfy:
    $\text{tl}(G)\leq\text{td}(G)\leq\text{ts}(G),$
    and
    $\text{tb}(G)\leq\frac{\text{ts}(G)}{2}.$
    [379]
  8. 8.
    The tree-stretch is bounded by the tree-breadth and tree-distortion as:
    $\text{ts}(G)\leq 2\cdot\text{tb}(G)\cdot\log_{2}n,$
    and
    $\text{ts}(G)\leq 2\cdot\text{td}(G)\cdot\log_{2}n.$
    [379]
  9. 9.
[…]
  10. 10.
[…]
  11. 11.
    $\text{Slimness}\leq\left\lfloor\dfrac{3}{2}\,\text{treelength}(G)\right\rfloor$ [906].
  12. 12.
    $\text{Thinness}\leq 3\,\text{treelength}(G)$ [906].
  13. 13.
    Branch-length is equivalent to Tree-length[1164].